OpenStudy (anonymous):

PELEASE HELP!! Find an equation for the nth term of the arithmetic sequence. -17, -14, -11, -8, ...

OpenStudy (anonymous):

that is not one of my answer choices: an = -17 + 3(n - 1) an = -17 x 3(n - 1) an = -17 + 3(n + 1) an = -17 + 3(n + 2)

OpenStudy (blank ):

It is the same as A. an = -17 + 3(n - 1)